Grade 12Physical Chemistry

K.E. of the electron is . Its de Broglie wave length is

Rahul Milind Surange
16 Years agoGrade 12

1 Answer

AskiitiansExpert Abhinav Batra
16 Years ago

Dear Rahul

we now from the de brogelie` law

p=h/λ

but p2/2m=K.E.

so p=√2mK.E.

so √2mK.E.=h/λ

hence λ=h/√2mK.E.=√(hπ/m) where m iS the maSS of electron

SubStituting the valueS h=6.626*10-34 m=9.1*10-31 π=22/7

λ=4.78*10-2 m

ALL THE best

AskIITIANsExpert

ABhinav Batra